# Linying Lv

### v3

@@ 3.2.2 Earnings Announcements
Building on research examining the interplay between analyst reports and corporate disclosures (Frankel et al., 2006; Chen et al., 2010; Huang et al., 2018), I analyze the information content of sell-side analyst reports surrounding earnings announcements.
Following Chen et al. (2010), I group analyst reports into 13 weekly bins around earnings announcement dates. Panel A in Table 5 presents the $R^{2}_{\mathrm{OOS}}$ and DM t-stats for weekly report samples. Consistent with Huang et al. (2018), analyst reports cluster predominantly in the immediate post-earnings announcement period. The information content only exhibits strong significance only in the first week, achieving an $R^{2}_{\mathrm{OOS}}$ of approximately 10% with a t-statistic of 10.17.
The variation in performance across weekly bins may reflect Ridge models capturing dominant patterns in analyst reports issued around earnings announcements. To address potential temporal variation, I train separate Ridge models on analyst reports released during earnings announcement periods and those issued beyond these periods, allowing the models to extract distinct information content based on report timing. Panel B in Table 5 compares the model performance within and beyond earnings announcement periods using 1-day, 2-day, 3-day, and 7-day windows. Models trained on earnings announcement periods consistently outperform those trained on non-announcement periods. The $R^{2}_{\mathrm{OOS}}$ and t-statistics exhibit a declining pattern as windows lengthen, decreasing from 11.84% to 7.29% over the earnings announcement week, suggesting markets particularly value swift analyst responses to earnings news.
To address concerns that analysts merely restate earnings call information and that market reactions primarily reflect earnings announcements rather than analyst insights, I conduct a comparative analysis of information content between these two sources. I focus on reports released one day after earnings announcements, when analysts would be most likely to simply repackage earnings information.
Column (1) of Panel C presents the $R^{2}_{\mathrm{OOS}}$ for $CAR_{[-1,+1]}$ using earnings conference call transcript embeddings as input. [^10] Due to the length of earnings conference call transcripts (median 13,029 tokens), I employ the LLaMA-3-8B model for text embeddings in Panel C, which accommodates 8,192 tokens versus the 4,096-token limit of the LLaMA-2-13B model used elsewhere. Texts are truncated at these limits to maximize information retention. The $R^{2}_{\mathrm{OOS}}$ reaches 4.20% with a DM t-statistic of 5.16. Since the conference call transcript data began in 2005, I retrained the Ridge regression model with analyst report text input to align the sample periods. Using report text embeddings from the day following earnings announcements, this analysis achieves an $R^{2}_{\mathrm{OOS}}$ of 9.72% (column 2), substantially outperforming conference call transcripts alone.
The combination of analyst report and earnings conference call embeddings yields further improvements, with the $R^{2}_{\mathrm{OOS}}$ increasing to 11.96% (column 3). A comparison t-statistic of 5.12 between transcript-only and combined models demonstrates analysts’ substantial incremental contribution. These findings collectively refute the notion that analysts merely restate earnings information, instead highlighting their distinct and timely value to the market.
@@ 3.2.3 What Content in Analyst Reports is Valued?
Having established the significance of report text information, I examine which specific content components most influence market response. I employ the Shapley value methodology to decompose the text-based $R^{2}_{\mathrm{OOS}}$ across the 17 topics detailed in Table LABEL:tab:_topic. The analysis uses sentence-segmented embeddings as report representations to mitigate cross-topic contextual effects. When switching from full-context to sentence-segmented embeddings, the $R^{2}_{\mathrm{OOS}}$ decreases by 2.25%, quantifying the magnitude of inter-sentence contextual information.
@@
Figure 2 illustrates the Shapley value decomposition, revealing a clear hierarchy in topic importance. Income Statement Analyses and Financial Ratio Analyses emerge as the most valuable components, contributing 67% and 45% of text-based $R^{2}_{\mathrm{OOS}}$, respectively. Investment Thesis and Valuation form a second tier of importance, with Shapley values of 1.76% and 1.51%. The remaining topics show minimal or negative Shapley values, indicating limited market reactions.
The findings’ robustness is supported by several additional analyses. Figure A4 shows that normalizing Shapley values by sentence count and token length maintains Income Statement Analyses’ dominance. Figure A5 demonstrates consistency in topic rankings across different industries.
The evolution of text-based $R^{2}_{\mathrm{OOS}}$ is depicted in Figure A6 of the Online Appendix. While a notable decline in $R^{2}_{\mathrm{OOS}}$ occurs between 2020 and 2021, likely reflecting COVID-19 pandemic disruptions in financial markets, the scaled Shapley value for Income Statement analysis (shown by the red line) exhibits remarkable stability. This persistence suggests that analyst commentary on income statements maintains its informational value even during market turbulence.
To disentangle whether topic importance stems from analysts’ expertise in certain areas (supply effect) or investors’ response (demand effect), Figure A7 compares Shapley values across analyst experience levels and brokerage firm sizes. The consistent topic hierarchy across all four subsamples (experienced/inexperienced analysts and large/small brokerages) indicates that topic rankings are not driven by variations in analyst expertise.
@@
I proceed to examine analysts’ dual roles in income statement analysis: their discovery role in collecting and generating otherwise unavailable information and their interpretation role in integrating and processing available information into more interpretable signals (Huang et al., 2018; Blankespoor et al., 2020). To facilitate the analysis, I categorize Income Statement sentences based on: (1) Information Acquisition versus Information Interpretation, and (2) Realized Income versus Expected Income. This categorization is achieved through a prompt-based classification system, as detailed in Prompt 3.
@@
Expected Income (1): Sentences containing subjective predictions, or expectations for future periods. Example: ‘The company expects to achieve revenue of \$550 million in fiscal year 2023 based on current market conditions.’
@@
Figure 3 provides a granular breakdown of $R^{2}_{\mathrm{OOS}}$ across 18 topics, including Income Statement subcategories. Income interpretation contributes three times as much to $R^{2}_{\mathrm{OOS}}$ as income acquisition, suggesting that markets value context and explanation over raw figures. The contribution from realized income discussions is triple that of income expectations, indicating markets place greater value on actual results than forward-looking projections. These patterns emphasizes that markets particularly value analysts’ ability to interpret realized income, beyond mere information reporting or future expectations.
These findings align with the results from section 3.2.2, where analyst reports released immediately after earnings announcements demonstrate the highest information content. Taken together, these results suggest that the market places the highest value on analysts who can rapidly digest earnings announcement information and provide insightful interpretations.
3.2.4 Comparison of Revision Reports and Reiteration Reports
Given the established relationship between market prices and analyst forecast revisions (e.g., Gleason and Lee, 2003; Asquith et al., 2005; Bradshaw et al., 2021), I examine the differential information content between revision and reiteration reports. Following the criteria of Huang et al. (2014), I classify reports as revisions when released within two days of the I/B/E/S forecast announcement date. This classification is applied independently across three forecast targets—recommendations, target prices, and earnings. Therefore, a single report may be classified as both an earnings forecast revision report and a recommendation reiteration report.
Table 6 presents out-of-sample model performance across reiteration and revision samples. The frequency of revisions varies considerably: earnings forecast revisions are most common (56.57% of reports), followed by target price revisions (31.10%), while recommendation revisions are relatively rare (2.29%). Recommendation revisions exhibit particularly strong explanatory power. Quantitative forecast revision measures achieve an $R^{2}_{\mathrm{OOS}}$ of 14.99% when accompanying recommendation revisions, while report text yields 16.80%. The combination enhances explanatory power to 22.63%. This finding complements the literature that attributes large stock price changes to analysts’ recommendation revisions (Loh and Stulz, 2011; Bradley et al., 2014).
Target price revisions, noted for their incremental information value due to the limitations of discrete recommendations (Bradshaw, 2002; Brav and Lehavy, 2003), show marked differences between revisions and reiterations. While reiteration reports demonstrate minimal information content (combined $R^{2}_{\mathrm{OOS}}$ of 3.58%), revision reports exhibit substantial value ($R^{2}_{\mathrm{OOS}}$ = 20.88%, t-stat = 11.18).
Earnings forecast revisions show a similar pattern. Reiteration reports yield insignificant $R^{2}_{\mathrm{OOS}}$, while revision reports demonstrate significant explanatory power. The combination of numerical and textual inputs for revision reports achieves an $R^{2}_{\mathrm{OOS}}$ of 16.37% (t-statistic = 11.42), highlighting revisions’ superior value over reiterations.
Overall, Table 6 shows that report content carries greater information content when accompanying forecast revisions rather than reiterations, with recommendation revision reports exhibiting the highest value.
3.2.5 Comparison with Analyst Report Sentiment
The sentiment analysis serves two objectives: to evaluate how LLMs perform compared to traditional NLP methods, and to assess the relative information content of embeddings versus tone measures. I compare tone measures derived from both Huang et al. (2014)’s Naive Bayes approach and BERT classification. I manually label 10,000 sentences as training data for both models. After training the models to categorize sentences as positive, neutral, or negative, I average these classifications at the report level to create $Tone\_NB$ and $Tone\_BERT$ measures.
Table 7 contrasts the performance of text tone measures against text embeddings. Panel A presents OLS regression results of three-day $CAR$ on tone measures. Column (1) shows a coefficient estimate of 0.022 for $Tone\_NB$, aligning with Huang et al. (2014)’s reported 0.0208. A one standard deviation increase in text tone favorableness corresponds to a 54 basis point increase in $CAR$, comparable to their reported 41 basis point increase in two-day abnormal returns. Column (2) reveals significantly positive coefficients for tone measures for both Income Statement and non-Income Statement topics. The results remain consistent under firm-fixed and year-fixed effects in column (3). BERT-based tone measures (columns 4-6) demonstrate substantially larger coefficient estimates with higher significance levels. A one standard deviation increase in BERT-based tone favorableness yields a 95 basis point increase in three-day $CAR$. When both measures are included (column 7), Naive Bayes-based tone coefficients become insignificant, indicating BERT’s superior sentiment capture.
Panel B of Table 7 compares the $R^{2}_{\mathrm{OOS}}$ across models. Tone inputs alone achieve $R^{2}_{\mathrm{OOS}}$ ranging from 0.05% to 3.68% in the first four rows. Including forecast revisions and text embeddings improves $R^{2}_{\mathrm{OOS}}$ to 9.49% and 12.27%, respectively. The panel’s final rows show that adding tone measures to text embedding models yields no statistically significant improvement in $R^{2}_{\mathrm{OOS}}$.
The results highlight two key advantages of LLMs over conventional NLP methods: stronger explanatory power for stock returns through better sentiment capture, and the ability to extract richer textual information beyond simple sentiment measures.
3.2.6 Alternative Machine Learning Models
I examine the information content using three alternative machine learning algorithms, following the same procedure of Ridge regressions. Results are presented in Table 8.
PLS regressions extract common components that correlate with the target variable. If text embeddings contain strong common components that correlates with stock returns, PLS should outperform. Using this method, I achieve an $R^{2}_{\mathrm{OOS}}$ of 8.26%, with year-to-year variations similar to Ridge models. DM tests confirm significance at the 1% level across all years.
The XGBoost algorithm, which builds sequential trees to combine weak estimations into stronger ones, allows for testing whether text embeddings contain discrete informational components of varying importance. Despite its ability to capture nonlinear relationships, XGBoost underperforms Ridge regressions with an $R^{2}_{\mathrm{OOS}}$ of 5.97% (DM t-stat = 6.76). This relatively weaker performance suggests that linguistic information is distributed across the embedding space rather than concentrated in specific dimensions.
Neural Network models, which underpin LLMs and transformer architectures, show superior performance. Given their interconnected node structure mirroring human neural networks, these models are well-suited to capture word relationships and contextual information. Using ensembles of five models for each configuration, Neural Networks achieve the highest $R^{2}_{\mathrm{OOS}}$ among tested algorithms, ranging from 10.42% to 12.12%. Performance improves with network depth, consistent with Kelly et al. (2022)’s virtue of complexity principle.
Two key findings emerge from these experiments. First, the significant $R^{2}_{\mathrm{OOS}}$ across models demonstrates meaningful information content in report text, particularly given the low signal-to-noise ratio in stock returns. Second, although Neural Networks achieve the highest performance, most of the textual information can be captured by simple linear models like Ridge regressions.
@@ 3.3 Information Value of Analyst Reports
To translate the statistical measure $R^{2}_{\mathrm{OOS}}$ into economic terms, this section quantifies the information value of analyst reports using the methodology established in Section 2.3. This measure captures the economic profit generated from reduced price uncertainty, while incorporating price impact considerations. The measure’s construction implies that both higher information content and lower price impact contribute positively to the estimated information value.
Table 9 presents the statistical evidence estimated with the delta method. For S&P 100 constituent stocks, sell-side analysts generate an average three-day window information value of \$0.47 million (SE = \$0.05 million). This value can be decomposed into its constituent components: textual content contributes \$0.38 million, while numerical revisions account for \$0.34 million. Extrapolating from the typical frequency of analyst coverage—approximately 15 report days per year—the annualized lower bound investors would expect to profit from early access to analyst reports on an S&P 100 firm amounts to \$6.89 million. The precision of these estimates is supported by consistently narrow standard errors and confidence intervals across different information value measures. The average price impact ($\widehat{\lambda}_{it}/P_{it_{-}}$) is 0.34, suggesting that a \$1 million order pushes the stock price up by 0.034%. The statistics demonstrate that analyst information value is economically meaningful, with both textual content and numerical forecasts contributing significantly.
Figure 4 illustrates the dynamics of analyst report information value. The analysis tracks both the strategic value of combined numerical and textual information (solid line) and the isolated value of textual information (dashed line). While both measures exhibit substantial temporal variation, they demonstrate an overall upward trend. The consistently higher combined value relative to textual value suggests complementarity between numerical and textual components.
The 95% confidence interval reveals increased variability in information value post-2020, characterized by wider confidence bounds and more pronounced peaks and troughs. The heightened volatility coincides with the COVID-19 pandemic period, which introduced substantial market uncertainty. The patterns demonstrate that the dynamic nature of analyst information value and its sensitivity to market conditions.
@@
The Online Appendix provides additional robustness checks. Table A7 presents alternative information value estimates using high-frequency realized return volatility measures and varied trade-signing algorithms, yielding estimates ranging from \$0.42 million to \$0.64 million. Figure A8 demonstrates that these alternative specifications produce consistent temporal patterns in information value evolution.

### v4

@@ 3.2.2 The Timing of Analyst Reports around Earnings Announcements
Building on prior research examining the interplay between analyst reports and corporate disclosures (Frankel et al., 2006; Chen et al., 2010), this section investigates the information content of analyst reports around earnings announcements.
Following Chen et al. (2010), I group analyst reports into 13 weekly bins relative to the earnings announcement date. Panel A in Table 5 presents the $R^{2}_{\mathrm{OOS}}$ and DM t-statistics for each bin. Report issuance peaks in the week immediately following earnings announcements. This is also the period when reports are most informative. The $R_{\mathrm{OOS}}^{2}$ reaches nearly $10\%$ (DM t-statistic $=10.17$ ), indicating significant market reaction to analysts’ timely interpretations of earnings news.
To account for the possibility that analyst information differs depending on its timing relative to an EA, I estimate separate Ridge models for reports issued within different event windows. This approach allows the models to learn distinct relationships between text and returns for reports published in information-rich versus information-sparse periods. Panel B of Table 5 compares the performance of models trained on reports from varying windows ($1,2,3$, and $7$ days post-EA). The models trained on reports issued most promptly after earnings consistently outperform those trained on other periods. Notably, the $R_{\mathrm{OOS}}^{2}$ declines from $11.84\%$ to $7.29\%$ as the training sample window expands from one to seven days, suggesting that the market places the highest value on immediate, timely analysis following corporate disclosures.
A critical question is whether analysts provide unique insights or merely act as conduits, reiterating information already disclosed in earnings conference calls. To test this, I conduct a direct comparison of the information content of analyst reports and the latest earnings conference call transcripts. [^8] Given the significant length of conference call transcripts (median 13,029 tokens), I employ the LLaMA-3-8B model for this analysis, as its 8,192-token context window is larger than the 4,096-token limit of the LLaMA-2-13B model used elsewhere. Texts are truncated at these limits. I deliberately focus on reports released one day after an EA, creating a “hard test” for the analyst’s contribution, as this is when both analyst reports and contemporaneous returns are more likely to be influenced by earnings conference call information. I estimate Ridge regressions that combine both information sources:
$CAR_{[-1,+1],it}=\beta_{0}+\beta^{\prime}y_{ijt}^{{\text{AI+EA}}}+\epsilon_{ijt},$
[…]
where $y_{ijt}^{\mathrm{AI}+\mathrm{EA}}$ is created by concatenating the embeddings from analyst report $j$ for stock $i$ on day $t$ with the embeddings from that stock’s most recent earnings conference call.
Panel C of Table 5 presents the results. A model using only the text embeddings of earnings conference call transcripts achieves an $R^{2}_{\mathrm{OOS}}$ of $4.20\%$ (Column 1). The transcript data starts in 2005. For a fair comparison, I retrain analyst report models over the same 2005-2023 sample period. This model achieves an $R_{\mathrm{OOS}}^{2}$ of $9.72\%$ (Column 2), substantially outperforming the explanatory power of the earnings conference call transcripts alone.
Finally, a model that combines embeddings from both analyst reports and earnings conference calls yields an $R_{\mathrm{OOS}}^{2}$ of $11.96\%$ (Column 3). The highly significant improvement over the transcript-only model (DM t-statistic = 5.12) demonstrates the substantial, incremental information content provided by analyst reports.
These findings collectively refute the notion that analysts merely restate earnings information. Instead, they emphasize analysts’ role in distilling and contextualizing corporate disclosures in ways that are uniquely informative to the market.
@@ 3.2.3 What Content in Analyst Reports is Valued?
Having established that analyst report text contains significant value-relevant information, this section investigates which specific content categories drive this informativeness. Using the Shapley value methodology, I decompose the text-based $R^{2}_{\mathrm{OOS}}$ across the 17 report topics pre-defined in Table LABEL:tab:_topic. To minimize the confounding effects of cross-topic attention, the analysis uses sentence-segmented embeddings rather than full-context embeddings. This adjustment results in a 2.25 percentage point drop in $R^{2}_{\mathrm{OOS}}$, a figure that quantifies the value of inter-sentence contextuality.
@@
Figure 2 illustrates the Shapley value decomposition, revealing a clear hierarchy in topic importance. Income Statement Analysis emerges as the single most valuable component. The Shapley value indicates it accounts for a relative contribution of 67% to the total $R^{2}_{\mathrm{OOS}}$, followed by Financial Ratio Analysis as the second most crucial topic. A second tier of importance includes Investment Thesis and Valuation, with Shapley values of 1.76% and 1.51%. The remaining topics exhibit minimal or even negative Shapley values, suggesting little contribution to reports’ market-moving power.
The robustness of this topic hierarchy is confirmed through several additional tests. Normalizing the Shapley values by sentence count and token length (Figure A4) confirms that the dominance of Income Statement Analysis is not merely an artifact of size effect. Furthermore, this ranking remains remarkably consistent across all Fama-French 12 industries (Figure A5). The Online Appendix (Figure A6) shows that while the absolute predictive power of text fluctuated over time, dipping during the COVID-19 pandemic, the relative contribution of Income Statement Analysis remained stable, underscoring its enduring information value.
To disentangle whether the topic importance stems from analyst expertise in certain areas (a supply-side effect) or from consistent investor demand for specific information (a demand-side effect), I examine how topic importance varies across different analyst and report characteristics. Figure A7 presents the Shapley value decompositions for six subsamples.
The topic hierarchy proves to be remarkably consistent across all comparison groups: experienced versus inexperienced analysts; analysts at large versus small brokerage firms; and reports that primarily use DCF-based versus multiple-based valuation models. This pervasive consistency suggests that the dominance of certain topics, such as Income Statement Analysis, does not merely reflect analysts’ differing abilities, professional experience, or chosen valuation methodologies. Rather, the evidence indicates that the hierarchy is driven by the market’s broad and persistent demand for these specific types of fundamental analysis.
@@
Given its paramount importance, I conduct a deeper analysis of the “Income Statement Analysis” topic. I examine analysts’ dual roles: a discovery role, where they generate or report otherwise unavailable information, and an interpretation role, where they process and contextualize available information into more digestible signals (Huang et al., 2018; Blankespoor et al., 2020). I sub-classify each sentence in this topic along two dimensions: (1) Information Acquisition vs. Interpretation and (2) Realized vs. Expected Income, using the classification framework detailed in Prompt 3.
@@
Expected Income (1): Sentences containing subjective predictions or expectations for future periods. Example: ‘The company expects to achieve revenue of \$550 million in fiscal year 2023 based on current market conditions.’
@@
Figure 3 presents the results of this granular decomposition. Analysts’ interpretation of income statement data accounts for nearly three times the explanatory power of their acquisition and reporting of raw data. This suggests that investors value analytical context and insight far more than mere financial numbers. Similarly, discussions of realized (historical) income contribute three times as much to the $R^{2}_{\mathrm{OOS}}$ as discussions of expected (future) income, indicating the market’s strong preference for analysis grounds in actual financial outcomes over forward-looking speculation.
These results reinforce the findings from Section 3.2.2 that the market reacts most strongly to reports issued immediately following earnings announcements. This section reveals that the most valued content within those reports is the interpretation of realized performance. Taken together, the evidence suggests that the primary value-add of sell-side analysts lies in their ability to rapidly digest complex, newly-released corporate disclosures and provide timely, insightful interpretations of realized income.
3.2.4 Comparison of Revision and Reiteration Reports
This section investigates the differential informativeness of analyst reports conditional on whether they contain forecast revisions or simple reiterations. While the market impact of forecast revisions is well-documented (e.g., Gleason and Lee, 2003; Asquith et al., 2005; Bradshaw et al., 2021), I test whether the overall information content of analyst reports is significantly greater in revision versus reiteration events.
Following the classification methodology of Huang et al. (2014), I identify a report as a revision report if it is released within two days that a new forecast is recorded in the I/B/E/S database. This classification is applied independently to recommendations, target prices, and earnings forecasts, acknowledging that a single report may contain a revision for one forecasting target (e.g., target price) while reiterating another (e.g., recommendation).
Table 6 presents the out-of-sample performance, estimated separately for revision and reiteration report subsamples. Descriptive statistics reveal that revisions are relatively common for earnings forecasts (56.6%) and target prices (31.1%) but are rare for stock recommendations (2.3%). The subsequent analysis demonstrates that the market response is overwhelmingly concentrated in revision reports.
Among the three forecasting targets, recommendation revisions are the most potent information events. For reports containing upgrades or downgrades, the combined numerical and textual information achieves an exceptionally high $R_{\mathrm{OOS}}^{2}$ of $22.63\%$. This finding aligns with prior literature documenting the significant market impact of recommendation changes (Loh and Stulz, 2011; Bradley et al., 2014).
Reports accompanied by target price revisions also show strong explanatory power. While reiteration reports produce a modest combined $R^{2}_{\mathrm{OOS}}$ of 3.58%, revision reports attain 20.88% (t-stat = 11.18). This stark contrast underscores the incremental information value of specific target price changes beyond the coarser recommendations (Bradshaw, 2002; Brav and Lehavy, 2003).
Earnings forecast revisions follow a similar pattern. Reiteration reports offer little explanatory power, whereas revision reports achieve a combined $R^{2}_{\mathrm{OOS}}$ of 16.37% (t-stat = 11.42), confirming that the market responds specifically to forecast revision events.
In summary, the evidence in Table 6 consistently demonstrates that the informational value of analyst reports is heavily concentrated in those that revise explicit quantitative forecasts. Among these, the rarest recommendation changes, appear to convey the strongest signal to the market.
3.2.5 Comparison of Analyst Report Sentiment Measures
This section conducts a sentiment analysis to serve two primary purposes: first, to benchmark the performance of an LLM-based sentiment classifier (BERT) against a Naive Bayes algorithm used in Huang et al. (2014); and second, to assess the incremental information content of high-dimensional text embeddings relative to a single sentiment dimension.
Specifically, I examine tone derived from a Naive Bayes classifier (${Tone}_{{NB}}$) and a BERT-based sentiment classifier (${Tone}_{{BERT}}$). Both models are trained on a manually labeled dataset of 10,000 sentences categorized as positive, neutral, or negative. I calculate report-level sentiment scores for each report as the number of positive sentences minus the number of negative sentences, scaled by the total number of sentences.
Panel A of Table 7 presents OLS regressions of the CAR on these tone measures. In Column (1), I replicate the original analysis of Huang et al. (2014). The coefficient on ${Tone}_{{NB}}$ is 0.022, closely matching their estimate of 0.021. A one-standard-deviation increase in this tone measure is associated with a 54 basis point increase in CAR, comparable to the 41 basis point impact they document over a two-day window. Column (2) adds granularity by separating tone into Income Statement and non-Income Statement content, both yielding statistically significant effects. Column (3) confirms the robustness of these findings with firm and year fixed effects.
Columns (4)-(6) show that the BERT-based tone measure has substantially stronger explanatory power. A one-standard-deviation increase in ${Tone}_{BERT}$ corresponds to a 95 basis point increase in CAR. In a direct “horse race” regression including both measures (Column 7), the coefficient on ${Tone}_{{BERT}}$ remains large and highly significant, while the coefficient on ${Tone}_{NB}$ becomes insignificant. This indicates that the more advanced, context-aware BERT model captures more value-relavant text sentiment than the Naive Bayes classifier.
Panel B of Table 7 compares the $R^{2}_{\mathrm{OOS}}$ across distinct inputs. Tone measures alone offer modest explanatory power, with $R^{2}_{\mathrm{OOS}}$ ranging from 0.05% to 3.68%. Combining tone with forecast revision measures increases performance to 9.49%. Further integrating text embeddings lifts it to 12.27%. Notably, adding tone measures to a model already incorporating text embeddings yields no significant incremental improvement.
Together, the findings highlight two key advantages of LLMs. First, for the specific task of classifying sentiment, LLMs like BERT more accurately capture nuanced tone and generate stronger signals than traditional NLP methods. Second, LLM embeddings are informationally superior to single sentiment scores. The fact that tone adds no value once embeddings are included suggests that the high-dimensional embeddings already capture the sentiment dimension, in addition to the complex, context-rich information that simple sentiment indicators miss.
@@ 3.3 Information Value of Analyst Reports
To translate the statistical measure of $R^{2}_{\mathrm{OOS}}$ into economic terms, I follow the framework of Kadan and Manela (2025). The framework develops an empirical measure for quantifying the value of insider information base on the theoretical model of Back et al. (2000).
I employ the measure for two primary reasons. First, the framework is rooted in a general setting of imperfect competition among informed traders, which closely aligns with the reality of multiple investors with asymmetric yet overlapping access to analyst information. Second, this measure provides a conservative lower bound for the value estimation, allowing quantification of the minimum potential value of analyst information without overstating its importance.
The economic value of information ($\widehat{\Omega}_{it}$) for stock $i$ on day $t$ is defined as the ratio of the explainable component of return variance to the standardized price impact of trading:
$\widehat{\Omega}_{it}=\frac{\text{Explainable Return Variance}}{\text{Standardized Price Impact}}=\frac{r_{it}^{2}-\left(r_{it}-\widehat{r}_{it}\right)^{2}}{\widehat{\lambda}_{it}/p_{it_{-}}}.$ (8)
Here, $\widehat{r}_{it}$ represents the estimated cumulative abnormal return $\widehat{CAR}_{[-1,+1]}$ for stock $i$ on day $t$, and $r_{it}$ denotes the realized $CAR_{[-1,+1]}$ (where $\hat{r}_{it}=\frac{1}{N}\sum_{j=1}^{N}\hat{r}_{ijt}$ if multiple reports exist). The numerator thus captures the portion of return variance explained by the analyst reports. The denominator consists of the estimated price impact, $\widehat{\lambda}_{it}$, scaled by the stock’s closing price at day $t-2$, $p_{it\_}$.
This measure provides an estimate of the potential trading profits available to strategic investors who can trade on the information contained in analyst reports. By construction, the value of information ($\widehat{\Omega}_{it}$) increases with the model’s explanatory power (a higher numerator) and decreases with market liquidity (a lower price impact, $\widehat{\lambda}_{it}$, in the denominator), linking statistical informativeness with economic relevance.
To estimate the price impact $\widehat{\lambda}_{it}$, I regress 1-minute log returns $r_{itk}$ on contemporaneous order flow $y_{itk}$:
$r_{itk}=\widehat{\lambda}_{it}y_{itk}+\epsilon_{it},$ (9)
where $r_{itk}=p_{it\tau_{k}}-p_{it\tau_{k-1}}$ is the log price change, and $y_{itk}=Y_{it\tau_{k}}-Y_{it\tau_{k-1}}$ denotes signed order flow over interval $[\tau_{k-1},\tau_{k}]$. $\widehat{\lambda}_{it}$ (Kyle’s lambda) represents how sensitive the price is to trading volume. Buy and sell trades are classified using the algorithm of Lee and Ready (1991). Robustness checks also consider methods from Ellis et al. (2000) and Chakrabarty et al. (2007). The three-day interval $T$ is partitioned into one-minute segments $\tau_{0},\tau_{1},\ldots,\tau_{K}$, yielding 1,170 observations (390 × 3) per regression.
Given that $\widehat{\Omega}_{it}$ is a ratio of two random variables measured with error, the interpretation of the information value estimate requires care. To address this challenge, I apply the delta method to estimate the mean and variance of the information value. The derivation of first-order estimations using the delta method is detailed in the Online Appendix E.
Table 8 reports the average analyst information value estimates. The average economic value of analyst reports for an S&P 100 stock is \$0.47 million over a three-day event window (SE = \$0.05 million). This figure represents the potential profits available to strategic investors who can trade on the combined numerical and textual information.
To understand the sources of this value, I estimate the information value derived from textual and numerical content separately. A model based solely on text embeddings yields information value of \$0.38 million, while a model based only on numerical forecast revisions yields \$0.34 million. The fact that the combined model’s value (\$0.47M) is less than the sum of the individual components suggests a significant overlap or interaction between these information sources. Extrapolating from an average of 15 report days per year, the annualized lower-bound estimate of investor profits from early access to analyst information is \$6.89 million for a typical S&P 100 stock.
A key input to this calculation is the price impact of trading. The average standardized price impact ($\widehat{\lambda}_{it}/p_{it\_}$) is 0.34, which implies that a \$1 million trade moves the stock’s price by approximately 3.4 basis points. The modest price impact of these highly liquid, large-cap stocks helps explain the substantial economic value.
Figure 4 plots the time series of information value estimates from 2015Q1 to 2023Q4. The solid line, representing the combined value from text and numbers, exhibits a mild upward trend, suggesting that the economic value of analyst reports has, on average, increased over the past decade. The persistent gap between the combined value and the value from text alone confirms the complementary nature of quantitative and qualitative information. Notably, the shaded 95% confidence band widens post-2020, signaling greater uncertainty and variation in information value during and after the COVID-19 pandemic. The fluctuations demonstrate the dynamic nature of analyst information value and its sensitivity to market conditions.
@@
Further robustness checks are detailed in the Online Appendix. Table A8 shows that using alternative high-frequency volatility measures and different trade-signing algorithms (e.g., Ellis et al., 2000; Chakrabarty et al., 2007) yields consistent estimates, with average information values ranging from \$0.42 million to \$0.64 million. Furthermore, Figure A8 demonstrates that the temporal trends in information value are consistent across these alternative specifications.